OpenStudy (anonymous):

Which of the following fractions would convert to a repeating decimal? a. 7/10 b. 3/20 c. 7/25 d. 2/13

OpenStudy (anonymous):

@texaschic101 can u help me?

OpenStudy (texaschic101):

(a) 7/10 = .7 (b) 3/20 = .15 (c) 7/25 = .28 (d) 2/13 = .15384615385 ==> this one

OpenStudy (anonymous):

thx, could u help me with more terminating and repeating decimals?

OpenStudy (texaschic101):

I am sorry but I really can't help right now. I have to go somewhere. I will be back on later if you still need help...again, I am sorry

OpenStudy (anonymous):

its ok, thanks anyway